# Alphazero-like Tree-Search can guide large language model decoding and training

### v1

@@ 4.2 Results and Discussion
To answer this question, we conduct initial experiments for one iterative update. We utilize MCTS-$\alpha$ with old policy $\pi_{\theta_{0}}$, value $v_{\theta_{0}}$ and ORM $\hat{r}_{\theta_{0}}$, to sample answers on the training dataset as an augmentation to the origin one. It will be further used to finetune these models to ($\pi_{\theta_{1}},v_{\theta_{1}},\hat{r}_{\theta_{1}}$). We include two baselines, RFT (Yuan et al., 2023b), which utilizes rejection sampling to finetune the policy, with different sampling numbers $k$ or top $N$, and PPO (Schulman et al., 2017) for the RLHF task. Note that PPO conducts multiple iterative updates. It is not fully comparable to our method and we only add it for completeness. Refer to Appendix D.7 and D.8 for more experimental details.
@@
In Table. 6, we list results of iterative update on the GSM8K and RLHF, covering greedy decoding and MCTS-$\alpha$ over all policy and value combinations. Our empirical results validate that TS-LLM can further train LLM policy, value and ORM, boosting performance with the new policy $\pi_{\theta_{1}}$, new value and ORM $\{v,\hat{r}\}_{\theta_{1}}$, or both $(\pi_{\theta_{1}},\{v,\hat{r}\}_{\theta_{1}})$ in CoT greedy decoding and MCTS-$\alpha$. $\pi_{\theta_{1}}$’s greedy performance is even slightly better than RFT which is specifically designed for GSM8k. We believe by further extending TS-LLM to multi-update, we can make it more competitive though currently $\pi_{\theta_{1}}$ still cannot beat PPO-based policy. For the training of value function and ORM (Question 3), we compare MCTS-$\alpha$ in Table 5. We train value and ORM in two paradigms, one ($\{v,\hat{r}\}_{\theta_{1}}$) is optimized from the initial weights and mixture of old and new tree-search data; another($\{v,\hat{r}\}_{\theta_{1}}^{\text{RL}}$) is optimized from $\{v,\hat{r}\}_{\theta_{0}}$ with only new tree-search data. This is called RL because training critic model in RL utilizes a similar process of continual training. The results show that $\{v,\hat{r}\}_{\theta_{1}}$ outperforms $\{v,\hat{r}\}_{\theta_{1}}^{\text{RL}}$ on both old and new policy when conducting tree search, contrary to the normal situation in traditional iterative RL training.
@@ Appendix A More Related Work and Comparisons
Recent efforts have been made to investigate non-linear reasoning structures such as trees(Jung et al., 2022; Zhu et al., 2022; Xu, 2023; Xie et al., 2023; Yao et al., 2023; Hao et al., 2023). Approaches for searching on trees with LLM’s self-evaluation have been applied to find better reasoning paths, e.g. beam search in Xie et al. (2023), depth-/breadth-first search in Yao et al. (2023) and Monte-Carlo Tree Search in (Hao et al., 2023). Compared with these methods, TS-LLM is a tree search guided LLM decoding framework with a learned value function, which is more generally applicable to reasoning tasks and other scenarios like RLHF alignment. TS-LLM includes comparisons between different tree search approaches, analysis of computation cost, and shows the possibility of improving both the language model and value function iteratively. The most relevant work, CoRe (Zhu et al., 2022), proposes to finetune both the reasoning step generator and learned verifier for solving math word problems using MCTS for reasoning decoding which is the most relevant work to ours. Compared with CoRe, in this work TS-LLM distinguishes itself by:
@@ Appendix B Limitation and future work
Appendix C Background of Monte Carlo Tree-search Algorihtms
Once we build the tree, we can use various search algorithms to find a high-reward trace. However, it’s not easy to balance between exploration and exploitation during the search process, especially when the tree is sufficiently deep. Therefore we adopt Monte Carlo Tree Search(MCTS) variants as choices for strategic and principled search. Instead of the four operations in traditional MCTS (Kocsis & Szepesvári, 2006; Coulom, 2006), we refer to the search process in AlphaZero (Silver et al., 2017a) and introduce 3 basic operations of a standard search simulation in it as follows, when searching actions from current state node $s_{0}$:
@@ Appendix C Background of Monte Carlo Tree-search Algorihtms
In this paper, we introduce 3 variants of MCTS based on the above basic operations.
Appendix D Experiment Details
D.1 Task setups
@@
RLHF We choose a synthetic RLHF dataset Dahoas  [^1] https://huggingface.co/datasets/Dahoas/synthetic-instruct-gptj-pairwise serving as the query data. We split the dataset to 30000/3000 as training and test set respectively. For the reward model, we choose reward-model-deberta-v3-large-v2222https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large-v2 from OpenAssistant, which is trained from several RLHF datasets.
D.2 SFT and value training details

### v2

@@ 4.2 Results and Discussions
We conduct initial experiments for one iterative update in GSM8k and RLHF alignment task. We utilize MCTS-$\alpha$ with old policy $\pi_{\theta_{0}}$, value $v_{\phi_{0}}$ and ORM $\hat{r}_{\phi_{0}}$, to sample answers on the training dataset as an augmentation to the origin one. It will be further used to finetune these models to ($\pi_{\theta_{1}},v_{\phi_{1}},\hat{r}_{\phi_{1}}$). We include two baselines, RFT (Yuan et al., 2023b), which utilizes rejection sampling to finetune the policy, with different sampling numbers $k$ or top $N$, and PPO (Schulman et al., 2017) for the RLHF task. Note that PPO conducts multiple iterative updates. It is not fully comparable to our method and we only add it for completeness. Refer to Appendix E.8 and E.9 for more experimental details and Appendix D.1 for an ablation about how to use the collected data to optimize the the value and ORM.
@@
In Table. 4, we list results of iterative update on the GSM8K and RLHF, covering greedy decoding and MCTS-$\alpha$ over all policy and value combinations. Our empirical results validate that TS-LLM can further train LLM policy, value and ORM, boosting performance with the new policy $\pi_{\theta_{1}}$, new value and ORM $\{v,\hat{r}\}_{\phi_{1}}$, or both $(\pi_{\theta_{1}},\{v,\hat{r}\}_{\phi_{1}})$ in CoT greedy decoding and MCTS-$\alpha$. $\pi_{\theta_{1}}$’s greedy performance is even slightly better than RFT which is specifically designed for GSM8k. We believe by further extending TS-LLM to multi-update, we can make it more competitive though currently $\pi_{\theta_{1}}$ still cannot beat PPO-based policy.
4.2.5 Ablation studies
[…]
Table 5: Path@1 results with different training settings.
Search Algorithms Training Setting Accuracy(%)
CoT-SC@10@ORM Pure, Less 55.5 $\pm$ 0.6
Pure 55.3 $\pm$ 0.5
Mixed 55.9 $\pm$ 0.7
BFS Pure, Less 50.0 $\pm$ 0.3
Pure 52.7 $\pm$ 0.8
Mixed 52.5 $\pm$ 1.3
MCTS-$\alpha$ Pure, Less 49.7 $\pm$ 1.1
Pure 52.7 $\pm$ 0.8
Mixed 51.9 $\pm$ 0.6
Search space and search width As discussed in Sec. 3.1, the search space is limited by maximum tree width. We demonstrate the influence introduced by different tree constructions on Game24 with different node expansion sizes. Specifically, we set the number of maximal expanded node sizes as 6, 20, and 50. Table 6 lists the Path@1 performance and the number of tokens generated comparing TS-LLM’s variants, CoT-SC and CoT-SC-Tree. The almost doubled performance boost from 43.8 to 80.7 indicates the impact of different expansion sizes and tree-width, improving TS-LLM’s performance upper bound. Refer to Appendix D.2 for additional results on GSM8K and ProntoQA.
Table 6: Game24 Path@1 result with different tree-width. Larger search space leads to much better performance.
Performance(%) / # tokens
Method width=6 width=20 width=50
MCTS-$\alpha$ 41.6 63.3 74.5
MCTS-Rollout 43.8 71.3 80.7
BFS-V 43.2 64.8 74.6
CoT-SC-Tree@10 38.8 48.3 48.3
CoT-SC@10 - - 52.9
@@ 5 Conclusion
6 Impact Statements
This paper presents work whose goal is to advance the field of Deep Learning. There are potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A More Related Work and Comparisons
Recent efforts have been made to investigate non-linear reasoning structures such as trees(Jung et al., 2022; Zhu et al., 2023; Xu, 2023; Xie et al., 2023; Yao et al., 2023; Hao et al., 2023). Approaches for searching on trees with LLM’s self-evaluation have been applied to find better reasoning paths, e.g. beam search in Xie et al. (2023), depth-/breadth-first search in Yao et al. (2023) and Monte-Carlo Tree Search in (Hao et al., 2023). Compared with these methods, TS-LLM is a tree search guided LLM decoding framework with a learned value function, which is more generally applicable to reasoning tasks and other scenarios like RLHF alignment. TS-LLM includes comparisons between different tree search approaches, analysis of computation cost, and shows the possibility of improving both the language model and value function iteratively. The most relevant work, CoRe (Zhu et al., 2023), proposes to finetune both the reasoning step generator and learned verifier for solving math word problems using MCTS for reasoning decoding which is the most relevant work to ours. Compared with CoRe, in this work TS-LLM distinguishes itself by:
@@ Appendix B Limitation and future work
Appendix C Backgrounds and details of each tree-search algorithms in TS-LLM
C.1 Preliminaries of Monte Carlo Tree-search Algorihtms
Once we build the tree, we can use various search algorithms to find a high-reward trace. However, it’s not easy to balance between exploration and exploitation during the search process, especially when the tree is sufficiently deep. Therefore we adopt Monte Carlo Tree Search(MCTS) variants as choices for strategic and principled search. Instead of the four operations in traditional MCTS (Kocsis & Szepesvári, 2006; Coulom, 2006), we refer to the search process in AlphaZero (Silver et al., 2017) and introduce 3 basic operations of a standard search simulation in it as follows, when searching actions from current state node $s_{0}$:
@@ C.1 Preliminaries of Monte Carlo Tree-search Algorihtms
C.2 Comparison of tree-search algorithms in TS-LLM
In this paper, we introduce three variants of MCTS based on the above basic operations. Among the 3 variants, MCTS-$\alpha$ is closer to AlphaZero(Silver et al., 2017) and MCTS is closer to traditional Monte-Carlo Tree Search(Kocsis & Szepesvári, 2006). While MCTS-Rollout is closer to best-first search or A*-like tree search.
The major difference between the first three MCTS variants and BFS-V/DFS-V adopted from the ToT paper(Yao et al., 2023) is that the first three MCTS variants will propagate (i.e. the backup operation) the value and visit history information through the search process. MCTS-$\alpha$ and MCTS-Rollout bacpropagate after the expand operation or visiting a terminal node. MCTS backpropagates the information only after visiting a terminal node. For DFS-V, the children of a non-leaf node is traversed in a non-decreasing order by value. For efficient exploration, we tried 2 heuristics to prune the subtrees, (1) drop the children nodes with low value by prune_ratio. (2) drop the children nodes lower than a prune_value.
Appendix D Extra Experiments and Discussions
D.1 Different Value Training of iterative update
Table 7: Different value training for iterative update on GSM8k
Method Policy Value Performance(%)
MCTS-$\alpha$ $\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\phi_{0}}$ 51.9 $\pm$ 0.6
$\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\phi_{1}}^{\text{RL}}$ 52.0 $\pm$ 0.5
$\pi_{\theta_{0}}$$\{v,\hat{r}\}_{\phi_{1}}$ 53.2 $\pm$ 0.3
MCTS-$\alpha$ $\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\phi_{0}}$ 54.1 $\pm$ 0.9
MCTS-$\alpha$ $\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\phi_{1}}^{\text{RL}}$ 55.2 $\pm$ 1.2
MCTS-$\alpha$ $\pi_{\theta_{1}}$$\{v,\hat{r}\}_{\phi_{1}}$ 56.5 $\pm$ 0.6
To figure out the best way of training the value function during iterative update. We also compare MCTS-$\alpha$ in Table 7. We train value and ORM in two paradigms, one ($\{v,\hat{r}\}_{\phi_{1}}$) is optimized from the initial weights and mixture of old and new tree-search data; another($\{v,\hat{r}\}_{\phi_{1}}^{\text{RL}}$) is optimized from $\{v,\hat{r}\}_{\phi_{0}}$ with only new tree-search data. This is called RL because training critic model in RL utilizes a similar process of continual training. The results show that $\{v,\hat{r}\}_{\phi_{1}}$ outperforms $\{v,\hat{r}\}_{\phi_{1}}^{\text{RL}}$ on both old and new policy when conducting tree search, contrary to the normal situation in traditional iterative RL training.
D.2 Results of different node expansion on tasks
[…]
Let us first clarify how we choose the specific tree max width. For tree-max-width $w$ in GSM8k, Game24, and ProntoQA, we first start with an initial value $w=6$. Then by increasing it (10 in GSM8K, 20 in Game 24, and 10 in ProntoQA), we can see the trends in performance and computation consumption. In ProntoQA/GSM8K, the performance gain is quite limited while the performance gain is quite large in Game24. So at last, we in turn try smaller tree-max-width in GSM8K/ProntoQA (3) and also try even larger tree-max-width (50) in Game24. Our final choice of $w$ (in Table 1) is based on the trade-off between the performance and the computation consumption. Currently, our selection is mainly based on empirical trials and it might be inefficient to determine the appropriate tree-max-width $w$. We think this procedure can be more efficient and automatic by comparing it with the results of CoT-SC on multiple samples to balance the tradeoff between performance and computation consumption. Because CoT-SC examples can already provide us with information about the model generation variation and diversity. We can also leverage task-specific features, e.g. in Game24, the correctness of early steps is very important, so a large $w$ can help to select more correct paths from the first layers on the search trees.
For the analysis of the results in Table 8, Table 9 and Table 10. We can mainly draw two conclusions aligned with that in Q1/Q2 in the main paper.
First, the overall trend that larger search space represents better tree-search performance still holds. For most tree-search settings, larger tree-max-width $w$ and search space bring in performance gain. The only exception happens at MCTS-Rollout on GSM8k decreases when the tree-max-width $w=10$, this is due to the limitation of computation(limitation on the number of generated tokens which is $51200$ per problem) is not enough in wider trees which results in more null answers. Despite the gain, the number of generated tokens also increases as the tree-max-width $w$ becomes larger.
Secondly, the conclusions in the main paper about comparing different search algorithms still hold. BFS performs pretty well in shallow search problems like (GSM8K/Game24). Though we can still see MCTS-$\alpha$ and MCTS-Rollout improve by searching in large tree-max-width (such as Game24 expanded by 50), the performance gain is mainly attributed to the extra token consumption and is quite limited. For deeper search problems like ProntoQA (15) and RLHF (64), the performance gap is obvious and more clear among all expansion widths. This aligns with our conclusion in Q1’s analysis.
Table 8: Path@1 metric on GSM8k with different node size.
Method Performance(%) / # tokens
expand by 3 expand by 6 expand by 10
MCTS-$\alpha$ 49.2 $\pm$ 0.04 460 51.9 $\pm$ 0.6 561 51.7 $\pm$ 0.5 824
MCTS-Rollout 47.2 $\pm$ 0.8 856 47.8 $\pm$ 0.8 3.4k 45.9 $\pm$ 0.9 7.1k
BFS-V 49.1 $\pm$ 0.8 260 52.5 $\pm$ 1.3 485 52.2 $\pm$ 0.9 778
[…]
Table 9: Path@1 metric on Game24 with different node size.
Method Performance(%) / # tokens
expand by 6 expand by 20 expand by 50
MCTS-$\alpha$ 41.6 $\pm$ 0.8 243 63.3 $\pm$ 1.9 412 74.5 $\pm$ 0.7 573
MCTS-Rollout 43.8 $\pm$ 5.3 401 71.3 $\pm$ 2.5 670 80.7 $\pm$ 1.5 833
BFS-V 43.2 $\pm$ 2.0 206 64.8 $\pm$ 2.9 370 74.6 $\pm$ 0.5 528
[…]
Table 10: Path@1 metric on ProntoQA with different node size.
Method Performance(%) / # tokens
expand by 3 expand by 6 expand by 10
MCTS-$\alpha$ 94.1 $\pm$ 0.1 151 99.4 $\pm$ 0.2 190 99.8 $\pm$ 0.2 225
MCTS-Rollout 85.9 $\pm$ 0.8 151 96.9 $\pm$ 0.6 210 99.3 $\pm$ 0.4 264
BFS-V 83.7 $\pm$ 1.0 105 94.4 $\pm$ 0.3 126 97.6 $\pm$ 0.3 145
[…]
D.3 Wall-time and Engineering challenges
Table 11, Table 12 and Table 13 show the wall-time of running different tree search algorithms implemented in TS-LLM , searching for one answer per problem (i.e. path@1). We also show the wall-time of CoT greedy decoding and CoT-SC@10 with ORM aggregation as comparisons. We record the wall-time of inferencing over the total test dataset of each task.
The experiments were conducted on the same machine with 8 NVIDIA A800 GPUs, the CPU information is Intel(R) Xeon(R) Platinum 8336C CPU @ 2.30GHz.
[…]
There are two more things we want to clarify. Firstly, as we mentioned in Appendix B, our current implementation only provides an algorithm prototype without specific engineering optimization. We find a lot of repeated computations are performed in our implementation when evaluating the child node’s value. Overall, there still exists great potential to accelerate the tree-search process, which will be discussed in the next paragraph. We are continuously working on this (we will discuss the engineering challenges in the next paragraph). Secondly, this wall-time is just Path@1 result so they have different token consumptions, and DFS-V, BFS-V, and MCTS will degenerate into greedy value search as we mentioned before. We are also working on monitoring the time consumption for Path@N results so we can compare them when given the same scale of token consumption.
Engineering challenges and potentials. Here we present several engineering challenges and potentials to increase the tree-search efficiency.
  • Policy and Value LLM with the shared decoder. Our current implementation utilizes separate policy and value decoders but a shared one might be a better choice for efficiency. If so, most extra computation brought by value evaluation can be reduced to simple MLP computation (from the additional value head) by reusing computation from LLM’s policy rollout. It can largely increase the efficiency. We only need to care about the LLM’s rollout computations under this setting.
  • KV cache and computation reuse. KV cache is used in most LLM’s inference processes such as the Huggingface transformer’s (Wolf et al., 2020) generation function. It saves compute resources by caching and reusing previously calculated key-value pairs in self-attention. In the tree search problem, when expanding or evaluating a node, all preceding calculations for its ancestor nodes can be KV-cached and reused. However, because of the large state space of tree nodes, we cannot cache all node calculations since the GPU memory is limited and the communication between GPU/CPU is also inefficient (if we choose to store such cache in CPU). More engineering work is needed to handle the memory and time tradeoff.
  • Large-batch vectorization. Currently, our node expansion and node evaluation are only vectorized and batched given one parent node. We may conduct batch inference over multiple parent nodes for large-batch vectorization when given enough computing resources.
  • Parallel tree-search over multi-GPUs. Our implementation handles each tree over a single GPU. AlphaZero (Silver et al., 2017) leverages parallel search over the tree(Segal, 2010), using multi-thread search to increase efficiency. In LLM generation setting, the main bottleneck comes from the LLM inference time on GPU. Thus more engineering work is needed for conducting parallel tree-search over multi-GPUs.
  • Tree-Search with speculative decoding Speculative decoding (Leviathan et al., 2023) is a pivotal technique to accelerate LLM inference by employing a smaller draft model to predict the target model’s outputs. During the speculative decoding, the small LLM gives a generation proposal while the large LLM is used to evaluate and determine whether to accept or reject the proposal. This is similar to the tree-search process with value function pruning the sub-tree. There exists potential that by leveraging small LLMs as the rollout policy while large LLMs as the value function, we can also have efficient tree-search implementations.
Table 11: Wall-time results on GSM8k
Method Overall Time(sec) Average Time(sec) #Average Token
CoT-Greedy 216.93 0.17 98
[…]
MCTS 378.13 0.29 486
MCTS-$\alpha$ 527.31 0.41 561
MCTS-Rollout 2945.94 2.27 3.4k
BFS-V 383.08 0.29 485
DFS-V 387.00 0.30 486
Table 12: Wall-time results on Game24
Method Overall Time(sec) Average Time(sec) #Average Token
CoT-Greedy 44.88 0.15 76
[…]
MCTS 81.22 0.27 371
MCTS-$\alpha$ 134.19 0.45 412
MCTS-Rollout 193.18 0.64 670
BFS-V 79.48 0.26 369
DFS-V 80.46 0.27 369
Table 13: Wall-time results on ProntoQA
Method Overall Time(sec) Average Time(sec) #Average Token
CoT-Greedy 74.09 0.15 77
[…]
MCTS 130.15 0.26 125
MCTS-$\alpha$ 236.26 0.47 190
MCTS-Rollout 238.62 0.48 210
BFS-V 130.35 0.26 126
DFS-V 130.47 0.26 126
D.4 Discussion about Shared LLM decoder for both policy and critic.
As we mentioned in Appendix D.3, using a shared decoder for the policy and value LLM might further improve the computation efficiency for the tree-search process. Therefore, we conducted an ablation to compare the model under the settings of a shared decoder and the setting of separated decoders on Game24.
We first describe the training setting of both types of models we compared. For the setting of separated decoder, we refer to Appendix E.2 for details about dataset and training hyperparameters. For the setting of shared decoder, we train the shared policy and value LLM with the same data used in the setting of separated decoder. During training, a batch from the supervised finetuning (SFT) dataset and a batch from the value training dataset are sampled, the total loss of the shared policy and value LLM is computed by $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{SFT}}+0.5\cdot\mathcal{L}_{\text{Value}}$, where $\mathcal{L}_{\text{SFT}}$ is the cross entropy loss of predicting the next token in the groundtruth answer and $\mathcal{L}_{\text{Value}}$ is the Mean Square Error loss as we described in Equation 1. The training is conducted on 8 NVIDIA A800 GPUs, using a cosine scheduler decaying from lr=2e-5 to 0.0 with a warmup ratio of 0.03, a batch size of 128 for the supervised finetuning dataset and a batch size of 128 for the value training dataset. And we trained the shared decoder model for 3 epochs. This training setting is the same as used in the separated decoder setting.
We will compare these two types of model in the perspective of performance and computation efficiency. All tree-search algorithms are conducted under the same hyperparameters as those in Table 3 on Game24 in which the tree-max-width $w$ is set to 20.
[…]
Next, we show some preliminary comparisons in Table 15 on the computation efficiency of the separated decoder model and the shared decoder model, from the results of expanding a tree node and evaluating its children. Specifically, Table 15 presents the node expansion time and value calculation time with/without KV Cache under token-level and sentence-level situations. For the token-level node, we set $w=50$ while for the sentence-level node, we set $w=20$. The results successfully present that a shared decoder can largely increase the computational efficiency for value estimation (20x in the token-level setting and 9x in the sentence-level setting).
In all, in this section, we initially conduct explorations on leveraging shared policy and value LLM decoder. The result proves the potential of computational efficiency for the shared structure. However, more work is needed to help the stability of policy/value performance.
Table 14: Comparisions of separated/shared LLM decoder policy and critic models on Game24
Method Performance(%) / # tokens
Separated Decoder Shared Decoder
CoT 12.7 76 16.3 166
[…]
MCTS-$\alpha$ 63.3 $\pm$ 1.9 412 64.1 $\pm$ 1.3 561
MCTS-Rollout 71.3 $\pm$ 2.5 670 70.6 $\pm$ 0.4 855
BFS-V 64.8 $\pm$ 2.9 370 63.0 $\pm$ 1.0 495
[…]
Table 15: Time (seconds) on policy expansion and value evaluation for a single tree node. When using a shared LLM decoder for policy and value LLM, we can use KV Cache for value calculation. It is much more efficient than a separate value decoder without KV cache.
Node Type Policy Expansion Value with Cache Value without Cache
Token-Level 0.067 0.074 2.02
Sentence-Level 0.165 0.122 1.03
Appendix E Experiment Details
E.1 Task setups
@@
RLHF We choose a synthetic RLHF dataset (Dahoas, ) [^1] https://huggingface.co/datasets/Dahoas/synthetic-instruct-gptj-pairwise serving as the query data. We split the dataset to 30000/3000 as training and test set respectively. For the reward model, we choose reward-model-deberta-v3-large-v2222https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large-v2 from OpenAssistant, which is trained from several RLHF datasets.
Chess Endgame Chess Endgame is introduced by Abdulhai et al. (2023). Chess endgames provide a simpler and more goaldirected variation of the chess task. A classic theoretical endgame position consists of a position where the only pieces on the board are the two kings and the queen. Although the board position appears simple, a sequence of carefully calculated moves is required to win. We followed the setting of Abdulhai et al. (2023) use an opponent of StockFish whose Elo is 1200. We modified the environment dynamics so that the agent fails as long as he makes an illegal move.
E.2 SFT and value training details